OpenStudy (ramennoodle01):

What is the x-intercept of f(x)=-(x+8)^2+12?

jimthompson5910 (jim_thompson5910):

Replace f(x) with 0. Then solve for x. \[\Large f(x)=-(x+8)^2+12\] \[\Large 0=-(x+8)^2+12\] \[\Large 0-12=-(x+8)^2+12-12\] \[\Large -12=-(x+8)^2\] \[\Large -(x+8)^2=-12\] \[\Large (x+8)^2=12\] Do you see what to do next?

OpenStudy (triciaal):

can you respond to the question?

OpenStudy (epistemophile):

Treat your f(x) as the value of y for which x is placed. You know that the x intercept occurs when y=0, so simply make f(x) valued at 0, and solve the other side of the equation

OpenStudy (mathmale):

You need to solve the equation (x+8)^2=12. There will be two results. Please check to ensure that one or both of these results is actually a solution of the original equation.